Find the length of RS if S is between R and T, the length of RS is one-third the length of RT , RS =3x -20 and ST = 4x-11.
Answer by cleomenius(959)

The line segment is RST.
Since RS is 1/3 of the total length, ST is 2/3.
Therefore, 2(RS) is = ST, and RS + ST = RT.
2(3x -20) = 4x - 11
6x - 40 = 4x - 11
2x = 29
x = 14.5
3(14.5) - 20 = 23.5
4x - 11 = 47
2(23.5) = 47
The length of RS is 23.5.
The length of ST is 47.
The total length of RT is 70.5.
Cleomenius.